# Simple Techniques for Enhancing Sentence Embeddings in Generative Language Models

### v1

@@ 4 Experiments
This section delineates the experimental results of our two novel prompting engineering strategies: Pretended CoT and Knowledge Enhancement. Initially, in Section 4.1, we evaluate the efficacy of Pretended CoT and Knowledge Enhancement applied alongside EOL on prevalent 7B-scale generative PLMs, including OPT, LLaMA, LLaMA2, and Mistral. Subsequently, in Section 4.2, we examine the influence of PLM parameter sizes on the performance of these techniques, utilizing OPT models spanning from 350M to 13B parameters.
@@ 4.1 Efficacy on 7B-Scale Generative PLMs
Aligning with the established experimental framework, we commence by loading the initial checkpoints of PLMs and directly assessing the Spearman correlation scores across seven STS benchmarks, eschewing any parameter updates. Throughout the experiments, the configurations of PromptEOL, Pretended CoT, and Knowledge Enhancement remain consistent with Table 4.
@@
Table 5 furnishes a comprehensive summary of the experimental findings. Notably, Pretended CoT and Knowledge Enhancement significantly amplify the models’ performances based on PromptEOL. This improvement is particularly remarkable within the contexts of LLaMA and LLaMA2, where the average Spearman correlation coefficient experiences an impressive uplift of 10.56% (escalating from 68.76 to 76.02) and 10.15% (advancing from 70.03 to 77.14), respectively. Given that these enhancements are attained merely by merely appending a fixed prefix to the input template, without the necessity for additional training data or external modifications, the effectiveness of these strategies is manifestly apparent. Furthermore, across all evaluated PLMs, the peak performance on the seven STS tasks is consistently realized through either Pretended CoT or Knowledge Enhancement, affirming the robustness and versatility of our approach.
@@
| Fine-tuning on Unsupervised Datasets | ||||||||
@@
| BERT avg. | 30.87 | 59.89 | 47.73 | 60.29 | 63.73 | 47.29 | 58.22 | 52.57 |
| BERT prompt | 60.96 | 73.83 | 62.18 | 71.54 | 68.68 | 70.60 | 67.16 | 67.85 |
| ST5-Enc | 34.97 | 60.19 | 47.59 | 66.40 | 70.62 | 62.83 | 63.57 | 58.02 |
@@
In addition, we compare our approach against three unsupervised fine-tuning approaches based on BERT model: BERT-flow, BERT-whitening [25], and SimCSE. By employing Knowledge Enhancement to refine LLaMA2’s raw embeddings, our strategy attains an average Spearman correlation score of 77.14 in direct inference settings, transcending the scores achieved by the aforementioned unsupervised schemes. Coupled with the GPU memory usage comparison illustrated in Table 1, it becomes apparent that our approach not only surpasses in terms of performance but also exhibits greater computational efficiency, highlighting its substantial promise for practical implementations.
@@ 4.2 PLM Performance across Different Scales
For PLMs of all sizes, optimal performance is invariably attained through either Pretended CoT or Knowledge Enhancement, reaffirming the scalability and adaptability of our strategy. Given that no adjustments are made to the manual templates as outlined in Table 4 based on model size disparities, this performance uplift showcases the exceptional plug-and-play nature of Pretended CoT and Knowledge Enhancement, underscoring their potential for seamless integration into new application scenarios with minimal additional effort.
The experimental outcomes further elucidate that Pretended CoT consistently delivers a more stable enhancement to the models’ raw embeddings compared to Knowledge Enhancement. It exhibits a universally positive impact across all PLMs evaluated in this study. On the contrary, Knowledge Enhancement does not uniformly surpass the baseline established by PromptEOL, particularly in smaller-scale models such as OPT. However, as model size increases to 13 billion parameters, Knowledge Enhancement demonstrates superior efficacy over Pretended CoT. This trend suggests that Knowledge Enhancement, which encompasses a broader token range and a more complex vocabulary, necessitates a higher level of textual comprehension from the model to effectively incorporate prior information, thus benefiting more from increased model capacity.
@@ 5 Analysis
This section delves into the underlying factors contributing to the success of Pretended CoT and Knowledge Enhancement strategies introduced in this study. We initiate our analysis by exploring whether these methodologies yield sentence embeddings with enhanced distribution characteristics. This examination, predicated on alignment and uniformity metrics which measure the semantic space properties of PLMs, is presented in Section 5.1. Subsequently, in Section 5.2, we employ a randomly selected sample to visualize the attention allocation during sentence representation computation, aiming to ascertain the effectiveness of Knowledge Enhancement in directing the model’s focus towards critical elements of the original sentence.
@@ 5.1 Alignment and Uniformity
With LLaMA2 as the backbone, we evaluate the average Spearman correlation scores alongside alignment and uniformity metrics for PromptEOL, Pretended CoT, and Knowledge Enhancement. These assessments involve utilizing the output vectors from either the PLM’s last or penultimate hidden layers as embeddings. The results are documented in Table 7.
@@ 5.2 Attention Allocation
The compelling empirical evidence presented in Section 4 has provided robust validation for the efficacy of this approach. To facilitate a more transparent and intuitive grasp of the working principle behind Knowledge Enhancement, we conduct an analysis on a randomly selected sentence (“a man is driving a car”) from the STS-B test set. Employing LLaMA2 as the base model, we compute the cosine similarity between the sentence embeddings produced by PromptEOL and Knowledge Enhancement, which are output vectors corresponding to the prompt’s final token, and the output vectors of each token in the original sentence. According to the self-attention computation mechanism of the Transformer [26] architecture , a higher similarity value indicates more focused attention on a given token.
@@ 6 Conclusion
In this study, we initially elucidate through a series of experiments that discussions surrounding the Explicit One-word Limitation should concentrate on the direct inference scenarios of generative Pre-trained Language Models (PLMs), as this technique is not inherently critical for discriminative models or for the fine-tuning of generative PLMs. Building on this foundation, we propose two innovative prompting engineering strategies to further enhance the quality of generative PLMs’ raw embeddings: Pretended Chain of Thought and Knowledge Enhancement. Our extensive experimentation confirms the efficacy and universality of these methods across a spectrum of models with varying types and parameter sizes. Remarkably, when paired with 7B-scale PLMs, our approach not only achieve performance on par with or even superior to unsupervised contrastive learning methods that have undergone fine-tuning but also demonstrate a reduction in GPU memory consumption. Considering the plug-and-play characteristics of Pretended Chain of Thought and Knowledge Enhancement, they possess a broad application prospect. Additionally, through a detailed examination of alignment and uniformity metrics alongside attention distribution analyses, we illuminate the contributory factors underpinning the success of these two schemes. Our code has been open-sourced for researchers in the related fields to verify and conduct further experiments.

### v2

@@ 4 Experiments
This chapter presents the experimental results of our two novel prompt engineering strategies: Pretended CoT and Knowledge Enhancement. Initially, Section 4.1 assesses the efficacy of these methods when integrated with Explicit One-word Limitation (EOL). This evaluation is conducted on prevalent 7B-scale generative PLMs, including OPT6.7b, LLaMA7b, LLaMA27b, and Mistral7b. Following this, Section 4.2 provides a comparative analysis of GPU memory consumption between fine-tuning and direct inference approaches. Finally, Section 4.3 examines the impact of PLM parameter sizes on the performance of these techniques, utilizing OPT models ranging from 350M to 13B parameters.
Regarding the experimental setup, we follow the established practices from prior research, assessing model performance across seven STS benchmarks via the SentEval [34] toolkit, with the Spearman correlation coefficient score as the core evaluation metric.
@@ 4.1 Efficacy on 7B-Scale Generative PLMs
Since our study primarily targets direct inference scenarios, we uniformly load the initial checkpoint of PLMs, eschewing any further parameter updates. In all experimental groups, the configurations of PromptEOL, Pretended CoT, and Knowledge Enhancement remain consistent with Table 3.
@@
Table 4 provides a comprehensive summary of the experimental findings. Notably, Pretended CoT and Knowledge Enhancement significantly amplify the models’ performances based on PromptEOL. This improvement is particularly remarkable within the contexts of LLaMA and LLaMA2, where the average Spearman correlation coefficient experiences an impressive uplift of 10.56% and 10.15%, respectively. Given that these enhancements are attained merely by appending a fixed prefix to the input template, without the necessity for additional training data or external modifications, the effectiveness of these strategies is manifestly apparent. Furthermore, across all evaluated PLMs, the peak performance on the seven STS tasks is consistently realized through either Pretended CoT or Knowledge Enhancement, affirming the robustness and versatility of our approach.
In addition, we compare our method against three unsupervised fine-tuning approaches based on BERT model: BERT-flow, BERT-whitening [25], and SimCSE. By employing Knowledge Enhancement to refine LLaMA27b’s raw embeddings, our strategy attains an average Spearman correlation score of 77.14 without any training, surpassing the results of the aforementioned unsupervised schemes.
@@
| Fine-tuned on Unsupervised Datasets | ||||||||
@@
| BERT avg.‡ | 30.87 | 59.89 | 47.73 | 60.29 | 63.73 | 47.29 | 58.22 | 52.57 |
| BERT prompt‡ | 60.96 | 73.83 | 62.18 | 71.54 | 68.68 | 70.60 | 67.16 | 67.85 |
| ST5-Enc§ | 34.97 | 60.19 | 47.59 | 66.40 | 70.62 | 62.83 | 63.57 | 58.02 |
@@
4.2 Computational Cost Comparison
This subsection quantifies the computational efficiency of direct inference methods utilizing raw embeddings from PLMs, in comparison to contrastive learning approaches that necessitate fine-tuning on extensive datasets. We document the GPU memory consumption for both approaches, with the results detailed in Table 5.
The results indicate that parameter updates on PLMs, particularly at the 7B-scale, require substantial computational power. In stark contrast, our Pretended CoT and Knowledge Enhancement strategies ingeniously leverage the inherent capabilities of generative PLMs without the need for any gradient updates, thereby reducing GPU memory usage to under 40GB. Given the exceptional performance demonstrated by Pretended CoT and Knowledge Enhancement in Section 4.1, they hold significant potential for practical implementations.
@@ 4.3 PLM Performance across Different Scales
For PLMs of all sizes, optimal performance is invariably attained through either Pretended CoT or Knowledge Enhancement, reaffirming the scalability and adaptability of our strategies. Given that no adjustments are made to the manual templates as outlined in Table 3 based on model size disparities, this performance uplift showcases the exceptional plug-and-play nature of Pretended CoT and Knowledge Enhancement, underscoring their potential for seamless integration into new application scenarios with minimal additional effort.
Furthermore, the experimental outcomes elucidate that Pretended CoT consistently delivers a more stable enhancement to the models’ raw embeddings compared to Knowledge Enhancement. It exhibits a universally positive impact across all PLMs evaluated in this study. On the contrary, Knowledge Enhancement does not uniformly surpass the baseline established by PromptEOL, particularly in smaller-scale models such as OPT350m. However, as model size increases to 13 billion parameters, Knowledge Enhancement demonstrates superior efficacy over Pretended CoT. This trend suggests that Knowledge Enhancement, which encompasses a broader token range and a more complex vocabulary, necessitates a higher level of textual comprehension from the model to effectively inject prior information, thus benefiting more from increased model capacity.
@@ 5 Analysis
This chapter delves into the underlying factors contributing to the success of Pretended CoT and Knowledge Enhancement strategies introduced in this study. We initiate our analysis by exploring whether these methodologies yield sentence embeddings with enhanced distribution characteristics. This examination, predicated on alignment and uniformity metrics that measure the semantic space properties of PLMs, is detailed in Section 5.1. Subsequently, in Section 5.2, we employ a randomly selected sample to visualize the attention allocation during sentence representation computation, aiming to ascertain the effectiveness of Knowledge Enhancement in directing the model’s focus towards critical elements of the original sentence.
@@ 5.1 Alignment and Uniformity
With LLaMA27b as the backbone, we evaluate the average Spearman correlation scores alongside alignment and uniformity metrics for PromptEOL, Pretended CoT, and Knowledge Enhancement. These assessments involve utilizing the output vectors from either the PLM’s last or penultimate hidden layers as embeddings. The results are documented in Table 7.
@@ 5.2 Attention Allocation
The compelling empirical evidence presented in Section 4 has provided robust validation for the efficacy of this approach. To facilitate a more transparent and intuitive grasp of the working principle behind Knowledge Enhancement, we conduct an analysis on a randomly selected sentence (“a man is driving a car”) from the STS-B test set. Employing LLaMA27b as the base model, we compute the cosine similarity between the sentence embeddings produced by PromptEOL and Knowledge Enhancement, which are output vectors corresponding to the prompt’s final token, and the output vectors of each token in the original sentence. According to the self-attention computation mechanism of the Transformer [26] architecture , a higher similarity value indicates more focused attention on a given token.
@@ 6 Conclusion
In this study, we initially elucidate through a series of experiments that discussions surrounding the Explicit One-word Limitation should concentrate on the direct inference scenarios of generative Pre-trained Language Models (PLMs), as this technique is not inherently critical for discriminative models or for the fine-tuning of generative PLMs. Building on this foundation, we propose two innovative prompt engineering strategies to further enhance the quality of generative PLMs’ raw embeddings: Pretended Chain of Thought and Knowledge Enhancement. Our extensive experimentation confirms the efficacy and universality of these methods across a spectrum of models with varying types and parameter sizes. Remarkably, when paired with 7B-scale PLMs, our approach not only achieve performance on par with or even superior to unsupervised contrastive learning methods that have undergone fine-tuning but also demonstrate a reduction in GPU memory consumption. Considering the plug-and-play characteristics of Pretended Chain of Thought and Knowledge Enhancement, they possess a broad application prospect. Additionally, through a detailed examination of alignment and uniformity metrics alongside attention distribution analyses, we illuminate the contributory factors underpinning the success of these two schemes. Our code has been open-sourced for researchers in the related fields to verify and conduct further experiments.